OpenStudy (anonymous):

Multiply these binomials (x - 9)(x + 9)

OpenStudy (anonymous):

=x^2+9x-9x-81 =x^2-81

OpenStudy (anonymous):

x(x+9)-9(x+9) x^2+29x+81 x^2+81

OpenStudy (radar):

These are the factors of the difference of two perfect square. Something you will soon recognize as you get more and more practice. The difference of two perfect squares in this case is \[x ^{2}-81\] which factors nicely to (x+9)(x-9) Putting it in the most general way \[a ^{2}-b ^{2}=(a+b)(a-b)\]